# rQUFguide.bib

### v1

@@ 3.2 Price Changes:
Price changes are one of the most important events in an LOB since it either signifies a trade that has happened or a change in the best quotes. To measure the behavior of mid-price changes in the Limit Order Book (LOB), we analyze the impact of trades on the LOB. Specifically, if the size of a trade (i.e., a market order) exceeds the instantaneous volume at the top of the book on either the bid or ask side, it leads to a queue depletion event. The remainder of the market order’s volume will then be filled from deeper levels of the LOB. The ratio of the trade size to the instantaneous top-of-the-book volume reflects the liquidity available at the top of the book. We consider two key statistics: the distribution of this ratio and the average amount (in price levels) by which a market order depletes the LOB. These statistics serve as our stylized facts for mid-price changes in the LOB. We show the mean trade induced price moves against varying relative tick sizes in Figure 3. On a log-log scale, again, we see a clear linear relationship between the two for small and medium tick stocks. There is some non-linearity for large tick stocks since the absolute returns are floored at half a cent. We show this non-linearity by depicting two regressions - one for large tick stocks (purple line) and one for all other stocks (red line). The slope of the fit for small/medium tick stocks is $-0.95$. Since relative tick-size here is proportional to the price of the security, we therefore conclude that the mean trade induced price is almost linearly related to the current price of the security. This experiment verifies the utility of using geometric scaling for price process models (the most simple one being the Geometric Brownian Motion).
Figure 23 demonstrates that as the relative tick size decreases, market orders deplete a larger proportion of the top quote size. This observation indicates that the top of the book in small-tick stocks is much more volatile compared to large-tick stocks. In Figure 4, we overlay the density of price moves resulting from a market order in the LOB. For large-tick stocks like SIRI, the probability of a price move exceeding one tick (e.g., one cent on the x-axis) is almost nonexistent. Conversely, for small-tick stocks, price moves of hundreds of ticks (e.g., one dollar) are relatively common. This discrepancy arises because the market order size relative to the top-of-the-book quotes is higher in small-tick stocks, combined with the fact that small-tick stocks typically have a sparser LOB. Sparse LOBs can lead to larger price moves due to queue depletion events.
Trade-induced mid-price moves are crucial for the price discovery process. The significant differences observed among various stocks in this stylized fact underscore the importance of accounting for such dynamics when choosing an LOB model. Additionally, it is important to note that the volatility process is scaled by the average magnitude of price changes. Models that do not account for multiple-tick price moves may underestimate the observed volatility.
3.3 Shape of the LOB:
In order to measure the distribution of available liquidity at various price points in a LOB, we count the total size of outstanding limit orders with respect to distance from the mid-price. This quantity, as a function of distance from mid-price, is what we define as the shape of the LOB. The shape of the Limit Order Book (LOB) plays a crucial role in understanding liquidity distribution. However, modeling deeper levels of the LOB increases both the dimensionality and complexity of the model, especially for small-tick stocks. In a generalized LOB model, it is essential to account for the vast variability in liquidity distribution across different relative tick sizes.
@@ 3.3 Shape of the LOB:
Figure 6: Depth with Maximum Liquidity of the LOB (log-log): Again we observe two regimes - the large tick names have the maximum liquidity at the top of the book however for small tick stocks, the maximum liquidity is correlated with the current rel. tick size as can be seen above. This correlation is power law with an exponent over rel. tick size of approximately $-1.5$.
Stock 25th Perc. 50th Perc. 75th Perc.
@@
Let the volume resting at a price level $p$ at time $t$ be denoted as $Q(p,t)$, which refers to the outstanding quotes at that price. The mid-price at time $t$ is denoted as $p_{\text{mid}}(t)$. We define the shape of the book as the time-weighted distribution of the volume $Q(p,t)$ as a function of the distance of the price level $p$ from the mid-price $p_{\text{mid}}(t)$. The shape of the book is calculated by first computing the volume at different price levels, weighted by the duration for which a particular shape persists. If a shape remains for 1 second, it is weighted 10 times more than one that persists for 0.1 seconds. The final distribution of volume for each day is referred to as the shape of the book for that day. Averaging over one calendar year gives the average shape of the book.
The shape of the book decays from its maximum at half a tick from the mid-price for large and medium-tick stocks, as illustrated in Figures 25(a) and 25(b). Notably, the shape profile decays more rapidly for large-tick stocks compared to medium-tick ones. For stocks like IBM, which are close to the boundary between medium and small-tick stocks, the maximum liquidity shifts deeper into the book. On the other hand, for small-tick stocks such as the one shown in Figure 25(c), the liquidity distribution is wider, with the maximum liquidity lying several ticks below the mid-price. Figure 5 further shows that within small-tick stocks, the shape becomes flatter as the relative tick size decreases.
Next, we calculate the maximum of the average shape of the book profile and show the results in Figure 6. As we can see, there are two regimes - large tick stocks have the maxima of the shape uniformly at half a tick away from the mid-price (fit shown in purple line). However we see an almost linear trend wherein the maximum liquidity shifts deeper in the book as we decrease the relative tick size (fit shown in red line). The slope of the red line is $-1.513$, therefore the relation is approximately power law with an exponent of $-1.5$ over the relative tick-size.
We also plot the empirical density of the instantaneous maxima to analyze the depth at which maximum liquidity lies in the LOB. The variability in the location of this maximum provides insights into how deep the liquidity is distributed in the book. The empirical density of the maxima of the liquidity as a function of distance from the mid-price, shown in Figure 26, indicates that the variance in the location of maximum liquidity increases as the relative tick size decreases. Even in small-tick stocks, there are significant instances where the maximum liquidity lies just half a tick from the mid-price. However, as depicted in Figure 21, the time-weighted likelihood of a one-tick wide spread is quite low.
@@
In Table 2, we display the various percentiles of the shape of the book for the 15 stocks tested. This statistical summary provides further insights into the variability of the liquidity distribution across different stocks and relative tick sizes.
In summary, the shape of the LOB is a stylized fact that must be carefully modeled, particularly in the context of small-tick stocks, where the liquidity distribution exhibits significant variability and complexity.
3.4 Sparsity of the LOB:
A notable feature in small-tick stocks’ LOB is that not every price level in the LOB will have liquidity. We instead observe a number of empty price levels in the LOB (Figure 10) between two successive quotes. This implies that market participants are comfortable with placing their orders deeper in the LOB to post a more passive price. In large-tick assets, the probability of execution (also known as fill probability in literature) drastically decreases if one places an order deeper in the book. Small-tick assets clearly do not follow this trend. We therefore measure the sparseness of the LOB’s quoted prices in the following.
(a) Empty Levels between Top and Top + 1 (log density)
(b) Variance of Volume Percentiles (log)
@@ 3.4 Sparsity of the LOB:
Figure 7: Sparsity in the LOB: (a) depicts that the distribution of the number of empty levels becomes flatter as we decrease the rel. tick size. Indeed we can clearly observe that for large tick stocks (dashed lines), the distribution is very narrow and most of the mass is at 0. However for small tick (dash-dotted lines), there is a significant probability of more than 10 empty levels in the LOB. In (b) we show the variability in the instantaneous shape of the LOB by calculating the variance of the percentiles of the shape of the LOB. This measure informs us about the variation of liquidity in the book and therefore is a proxy measure for sparsity. As we can see in this figure, there is continous trend of increasing variance as we decrease the rel. tick size for all percentiles of the shape of the LOB.
Figure 8: Instantaneous vs Average Shape of the LOB: This measure indicates how different, on a time-weighted average scale, is the instantaneous shape of the LOB from the average shape. This difference indicates whether the shape of the book remains persistent in the LOB or does the LOB’s liquidity vary a lot owing to the sparse levels in the LOB. As we can see clearly here, there is an increasing trend as we decrease the relative tick size however this relation seems to be non-linear.
@@
We are interested in the number of empty levels between two quotes on either the ask or bid side of the book. Mathematically, let $n_{e}(p,t)$ denote the number of empty levels between two quotes at price level $p$ at time $t$. We calculate $n_{e}$ for the top ten quotes in the LOB on both a time-weighted daily average basis and an instantaneous basis. The time-series of the time-weighted daily mean, along with the empirical distribution (log-scaled) of the instantaneous number of empty levels, reveals a trend: as the relative tick size decreases, the distribution becomes flatter and heavier-tailed. This trend is evident in Figure 7(a), where the number of empty levels between the best and second-best quotes increases with decreasing relative tick size. Figure 22 further confirms that this pattern persists for deeper levels in the book.
Next, we calculate the percentiles of the shape of the book at the 10th, 25th, 75th, and 90th percentiles. The conjecture is that these percentiles will remain relatively invariant if the book is dense, meaning there are close to zero empty levels in the LOB. Thus, the variance of these percentile levels serves as a proxy measurement of the LOB’s sparsity. As shown in Figure 7(b), there is a steady relationship between the variance of the levels where certain percentages of total liquidity can be found and the relative tick size. This variance increases with decreasing tick size, indicating greater sparsity.
Finally, to measure sparsity robustly, especially against small orders lying between two price levels with significant quote sizes, we compute the Wasserstein Distance (or Earth-Mover’s Distance) between the average shape of the book and the instantaneous shape of the book. This metric helps differentiate between stocks where the average shape consists of sparse Dirac probability masses and those where both the average and instantaneous shapes are dense. Figure 8 illustrates a continuous increase in the Earth-Mover’s Distance with decreasing relative tick size, suggesting that the dissimilarity between average and instantaneous shapes increases as tick size decreases.
In summary, Figures 7(a) and 22, along with the calculations of percentiles and Wasserstein Distance, collectively indicate that the sparsity of the LOB increases as the relative tick size decreases. This finding underscores the challenge of accounting for sparsity in LOB models, as modeling order arrivals across numerous price levels, with potential correlations among them, presents a high-dimensional problem further complicated by lower signal-to-noise ratios at deeper levels in the book.
@@
3.5 LOB Events and their Endogenous Excitation:
@@
As previously discussed, a Limit Order Book can be modeled as a queuing system, where various types of events contribute to the formation and evolution of these queues. Empirical observations suggest that the realizations of these events exhibit both cross-correlation with other events and autocorrelation with themselves. This phenomenon is commonly referred to as market endogeneity ([bouchaud2018trades]), reflecting the idea that market activity is influenced by its own past behavior. Our focus lies in analyzing the conditional probability distribution of these events, given the history of prior events, in order to capture the feedback mechanisms inherent in the LOB dynamics.
We study how the occurrence of certain events in the Limit Order Book (LOB) affects the likelihood of other events happening shortly afterward. Specifically, we look at the probability of an event of type $e^{(1)}$ occurring within the next second after an event $e^{(2)}$ has just happened. This helps us understand how one type of event can ”trigger” or increase the chances of another type of event. We express this by calculating the ratio of this conditional probability (i.e., the chance of $e^{(1)}$ given $e^{(2)}$) to the overall probability of $e^{(1)}$ occurring in one second. This ratio is known as Conditional Probability Leverage. We perform this analysis for three types events : Limit Orders (LO), Cancel Orders (CO), and Market Orders (MO), and extend it to all events affecting the top 10 levels of the book.
Additionally, inspired by [LuAbergel2018], we calculate the unconditional probability of various events (like Limit Orders, Immediate or Cancel Orders, etc.) happening at different distances from the mid-price. We denote these events by $e_{x}$, where $x$ is positive for events on the ask side and negative for events on the bid side. We then determine how likely it is for an event $e^{(1)}_{x}$ to occur given that a different event $e^{(2)}_{y}$ happened just before. We use this information to compute a measure called Leverage, which is the ratio of the conditional probability of $e^{(1)}_{x}$ following $e^{(2)}_{y}$ to the unconditional probability of $e^{(1)}_{x}$ i.e. $\frac{P(e^{(1)}_{x}|e^{(2)}_{y})}{P(e^{(1)}_{x})}$. We plot these results on a log scale for better clarity.
Figure 9 (and further Figure 24) show that the patterns in these leverage plots are similar across different tick sizes. This means that the way events influence each other is consistent, regardless of how large or small the tick size is. These findings provide insights into how the structure of the order book behaves and helps us understand the dynamics of event interactions within the LOB. Here we enlist some more observations from these plots:
@@
Figure 9: Leverage (from top): Darker the hue, more the leverage - The left hand side subplot shows the unconditional probability of the events while the leverage plot shows the conditional probability ratio. Note the similarities between the two figures (and further in Figure 24). This stylized fact motivates building a generalized self and cross excitory model for the 4 types of events shown here across all types of stocks, albeit with different scaling.
@@
    Limit Orders at the top of the book behave remarkably different to the deeper levels.
@@
    Cancels and Limit Orders behave very similar to each other.
@@
    Unsurprisingly, the unconditional probability of In-Spread Limit Orders (above 1 tick distance from the top) is very small for large and medium tick stocks.
@@
    In-Spread Limit Orders are both self excited and cross excited by opposite end LOs and COs.
@@
    In-Spread Limit Orders excite other events very differently between small and large tick stocks.
  6. 6.
@@
    Market Orders only occur at the top of the book, however we can see very similar patterns across all types of stocks for Market Orders.
We therefore conclude that in-spread orders should be modelled differently for different types of stocks however there is uniformity in other order types’ behaviour. We observe that our choice of modelling the top of the book as a separate entity to the deeper part of the book has some support.
4 A Simple Model

### v2

@@ 2.2 Mid-Price Changes: Probability Distribution vs Relative Tick-Size
In Figure 4, we overlay the density of price moves resulting from a market order in the LOB. For large-tick stocks like SIRI, the probability of a price move exceeding one tick (e.g., one cent on the x-axis) is almost zero. Conversely, for small-tick stocks, price moves of hundreds of ticks (i.e. one dollar) are relatively common. This discrepancy arises because the market order size relative to the top-of-the-book quotes is higher in small-tick stocks (as shown in the Appendix in Figure 21), combined with the fact that small-tick stocks typically have a sparser LOB (discussion continued in Section 2.4). Sparse LOBs can lead to larger price moves due to queue depletion events.
2.3 Shape of the LOB: Average Volume Distribution and its Maxima vs Relative Tick-Size
A passive liquidity seeker faces a trade-off when posting a limit order in the LOB: they can either post at the best quote for a higher fill probability or at deeper levels to save a few cents, sacrificing fill likelihood. This trade-off is informed by the distribution of available volume in the LOB across price levels. When normalized to sum up to one, we define this volume distribution as the shape of the LOB. Naturally, we expect the most liquidity to be available at the best bid or ask since they have higher fill rates. While this assumption seems to be true for large tick stocks, breaks down in medium or small tick stocks as we show in the following. This implies that the maximum liquidity of the LOB actually lies at deeper price levels and that the top of the book is not necessarily the best place to post a passive limit order for these categories of stocks.
@@ 2.3 Shape of the LOB: Average Volume Distribution and its Maxima vs Relative Tick-Size
Stock 1st Quartile 2nd Quartile 3rd Quartile
@@
In order to measure the distribution of available liquidity at various price points in a LOB, we count the total size of outstanding limit orders with respect to distance from the mid-price. This quantity, as a function of distance from mid-price and normalized to sum to one, is what we define as the shape of the LOB. Let the volume resting at a price level $p$ at time $t$ be denoted as $Q(p,t)$, which refers to the outstanding quotes at that price. We define the shape of the book as the time-weighted normalized distribution of the volume $Q(p,t)$ as a function of the distance of the price level $p$ from the mid-price $p_{\text{mid}}(t)$ and denote it by $\overline{q}(x)$ where $x:=\frac{p-p_{\text{mid}}(t)}{\delta}$ is the distance from mid-price in ticks. Similar to Equation 3, we have for index $i=1,...,N$ and time $0\leq t_{i}\leq T$, LOB volume $Q(p,t_{i})$ lasting for $\Delta t_{i}$ seconds, so the calculation of $\overline{q}(x)$
$\overline{q}(x):=\frac{\sum^{N}_{i=1}Q(p_{\text{mid}}(t)+\delta x,t_{i})\Delta t_{i}}{T}$ (6)
Figure 6: Depth with Maximum Liquidity of the LOB (log-log): Again we observe two regimes - the large tick names have the maximum liquidity at the top of the book however for small tick stocks, the maximum liquidity is correlated with the current rel. tick size as can be seen above. This correlation is power law with an exponent over rel. tick size of approximately $-\frac{3}{2}$.
@@
We overlay the average shape of the book for all 15 stocks in Figure 5. Figure 5 further shows that within small-tick stocks, the shape becomes flatter as the relative tick size decreases. In Table 2, we display the various quartiles of the shape of the book for the 15 stocks tested. This statistical summary provides further insights into the variability of the liquidity distribution across different stocks and relative tick sizes. Further in the Appendix, in Figure 22, we show some per stock daily and averaged over the year shapes of the LOB for a large, medium and small tick asset.
Next, in Figure 6, we calculate the maximum of the average shape of the book profile (i.e. $\text{arg max}_{x}\overline{q}(x)$) and plot the results against relative tick-size. As we can see, there are two regimes - large tick stocks have the maxima of the shape uniformly at half a tick away from the mid-price (fit shown in purple line). However we see an almost linear trend wherein the maximum liquidity shifts deeper in the book as we decrease the relative tick size (fit shown in red line). The slope of the red line is $-1.513$, therefore the relation is approximately power law with an exponent of $-\frac{3}{2}$ over the relative tick-size:
$\text{arg max}_{x}\overline{q}(x)\propto\epsilon^{-\frac{3}{2}}$ (7)
[bouchaud2018trades] mention the same fact that the most liquidity in small tick stocks actually lies much deeper in the LOB however to the best of our knowledge, our work is the first to form a statistical relation between this proxy of ‘market depth’ and relative tick-size or mid-price. Since $x$ is measured from the mid-price, and since the spread is scaling with the exponent of $-1$, the maximum volume’s distance from the respective top should scale in a square-root manner with the inverse of relative tick-size from our observations. In the Appendix, we also plot the empirical density of the instantaneous maxima to analyze the depth at which maximum liquidity lies in the LOB in Figure 23.
2.4 Sparsity of the LOB: Empty Price Levels and their effect on the instantaneous LOB’s Shape
@@
Figure 7: Empty Levels between Top and Top + 1 (log density) : This depicts that the distribution of the number of empty levels becomes flatter as we decrease the rel. tick size. Indeed we can clearly observe that for large tick stocks (dashed lines), the distribution is very narrow and most of the mass is at 0. However for small tick (dash-dotted lines), there is a significant probability of more than 10 empty levels in the LOB.
@@
A notable feature in small-tick stocks’ LOB is that not every price level in the LOB will have liquidity. We instead observe a number of empty price levels in the LOB (an illustration of an LOB snapshot is provided in Figure 10) between two successive quotes. As mentioned previously, the probability of execution (also known as fill probability in literature) drastically decreases if one places an order deeper in the book. However in the following we show that the book becomes more and more sparse as we decrease the relative tick-size. This implies that market participants are comfortable with placing their orders deeper in the LOB to post a more passive price in small tick stocks. We are interested in the number of empty levels between two quotes on either the ask or bid side of the book. Mathematically, let $n_{e}(i,t)$ denote the number of empty levels between two successive quotes, which have ranks $i$ and $i+1$, at time $t$. For example, $n_{e}(1,t)$ denotes the number of empty levels between the best quote and the second best quote and $n_{e}(2,t)$ is the same between second best and third best quotes. We calculate $n_{e}(i,t)$ for the top ten quotes in the LOB on both a time-weighted daily average basis and an instantaneous basis. We plot the empirical distribution (log-scaled) of $n_{e}(i,t)$ for $i=1$ in Figure 7. Notice that as the relative tick size decreases, the distribution becomes flatter and heavier-tailed. Further in the Appendix, in Figure 24, we confirm that this pattern persists for deeper levels in the book.
Figure 8: Typical vs Average Shape of the LOB: This measure indicates how different, on a time-weighted average scale, is the instantaneous shape of the LOB from the average shape. This difference indicates whether the shape of the book remains persistent in the LOB or does the LOB’s liquidity vary a lot owing to the sparse levels in the LOB. As we can see clearly here, there is an increasing trend as we decrease the relative tick size however this relation seems to be non-linear. The error bars denote the variance of the Wasserstein Distance observed in our dataset.
Finally, in Figure 8, to measure sparsity robustly, especially against small orders lying between two price levels with significant quote sizes, we compute the Wasserstein Distance (or Earth-Mover’s Distance) between the average shape of the book $\overline{q}(x)$ and the ‘typical’ shape of the book $q(x,t)$ at time $t$. The typical shape of the LOB is defined as the normalized distribution of volume as a function of distance from mid-price in ticks i.e.
$q(x,t):=\frac{Q(p_{mid}(t)+\delta x,t)}{\sum_{y}Q(p_{mid}(t)+\delta y,t)}$ (8)
This metric helps differentiate between stocks where the average shape consists of sparse ‘Dirac’ probability masses and those where both the average and typical shapes are dense. For two discrete PDFs $u(x),v(x)$ and their respective CDFs $U(x),V(x)$, discrete Wasserstein Distance $l(u,v)$ is defined as:
$l(u,v)$ $:=\sum_{x}|U(x)-V(x)|\Delta x$ (9)
We calculate $l(\overline{q}(.),q(.,t_{i}))$ for all timestamps $t_{i}$ and calculate the time-weighted average which we denote by $\overline{l}(\overline{q},q)$, dropping the time element of $q(.,.)$ by slight abuse of notation, and time-weighted variance i.e. $Var(l(\overline{q},q)):=\frac{\sum^{N}_{i=1}(l(\overline{q}(.),q(.,t_{i}))-\overline{l}(\overline{q},q))^{2}\Delta t_{i}}{T}$ of this quantity. Figure 8 therefore illustrates a continuous increase in the Earth-Mover’s Distance with decreasing relative tick size, suggesting that the dissimilarity between average and typical shapes increases as tick size decreases.
@@
2.5 LOB Events and their Endogenous Excitation: Universality across Relative Tick-Sizes
@@
As previously discussed, a Limit Order Book can be modeled as a queuing system, where various types of events contribute to the formation and evolution of these queues. Empirical observations suggest that the realizations of these events persistently exhibit both cross-correlation with other events and autocorrelation with themselves. This phenomenon is commonly referred to as market endogeneity ([bouchaud2018trades]), reflecting the idea that market activity is influenced by its own past behavior. Our focus lies in analyzing the conditional probability distribution of these events, given the history of prior events, in order to capture the feedback mechanisms inherent in the dynamics of the LOB.
Inspired by [LuAbergel2018], we study how the occurrence of certain events in the LOB affects the likelihood of other events happening shortly afterward. Specifically, we look at the empirical probability of an event happening at different distances from the respective top of the book (in ticks) of type $e^{(2)}_{y}$, where $y$ is positive for events on the ask side and negative for events on the bid side, occurring given that an event $e^{(1)}_{x}$ has just happened. This helps us understand how one type of event can “trigger” or increase the chances of another type of event. We express this by calculating the ratio of this conditional probability (i.e., the chance of $e^{(2)}_{y}$ given $e^{(1)}_{x}$) to the unconditional probability of $e^{(2)}_{y}$ occurring i.e. $\frac{P(e^{(2)}_{y}|e^{(1)}_{x})}{P(e^{(2)}_{y})}$. This ratio is known as Conditional Probability Leverage or simply ‘Leverage’. We perform this analysis for four types of events : Passive Limit Orders (LO) i.e. LOs which do not change the mid-price, Active or ‘In-Spread’ Limit Orders (IS) i.e. LOs which are posted at an improved price than the respective top price, Cancel Orders (CO), and Market Orders (MO), and extend it to all events affecting the top 10 levels of the book. We note that while there exist several more order types offered by the exchanges (for eg. fill-or-kill, stop loss etc), these four types of events are the most basic ones. These exotic orders can be thought of, mathematically, as a combination of these four basic order types. To facilitate a qualitative analysis, we plot the leverage for all these four events on a 2D grid with the x-axis representing the previous event i.e. $e^{(1)}_{x}$ in the example above and the y-axis the future event i.e. $e^{(2)}_{y}$. We plot these results on a log-log scale over $x,y$ for better clarity. As can be seen in the colormaps of these figures, a red shade corresponds to a conditional probability leverage higher than one which corresponds to an excitation by the preceding event to the next event. Similarly a blue shade corresponds to inhibition and white shade corresponds to independence between the two events.
@@
Figure 9: Leverage (from top of the book): Darker the hue, more the leverage - X-Axis is the preceding event and the Y-axis is the future event. Therefore one should read, for example, the value of the Figure 9(a) at $(x,y)=(LO_{0},LO_{-9})$ as equal to $\frac{P(LO_{0}|LO_{-9})}{P(LO_{0})}$. Note the similarities between the two figures (and further in Figure 26). This stylized fact motivates building a generalized self and cross excitory model for the 4 types of events shown here across all types of stocks, albeit with different scaling.
In Figure 9 (and further in the Appendix in Figure 26), we show that the patterns in these leverage plots are similar across different tick sizes, albeit with different scaling. In this figure we make use of a simple linear-log scale where we show the first 10 price queues (i.e. 10 ticks away from the respective top) for all kinds of stocks. From there onwards until the 95th percentile of the shape of the LOB, we use a log-scaling. This scaling is chosen in this way to enforce a higher resolution near the top of the LOB in these plots and to facilitate the aggregation of multiple price levels of the deeper parts of the LOB.
This uniformity in structure implies that the way events influence each other is consistent, upto a linear-log scaling, regardless of how large or small the tick size is. These findings provide insights into how the structure of the order book behaves and helps us understand the dynamics of event interactions within the LOB. Here we enlist some more observations from these plots:
@@
    Limit Orders at the top of the book behave remarkably different to the deeper levels: For example, in Figure 9(a), $x=LO_{0}$ mostly inhibits other orders (i.e. the tiles are blue colored) however $x=LO_{2}$ excites $y=LO_{0/1/2}$. A similar trend is observed in Figure 9(b).
@@
    Cancels and Limit Orders behave very similar to each other: This is except for the off-diagonal terms we see in $LO\rightarrow CO$ and vice-versa. The off-diagonal terms implies that deeper LOs lead to cancellations nearer to the top. Additionally deeper COs lead to LOs nearer to the top as well. Interestingly, this behaviour disappears for Figure 9(b). However we note that this transition is smooth from large to medium to small tick stocks (Figure 26(a) in the Appendix has both diagonal and off-diagonal terms).
@@
    In-Spread Limit Orders are both self excited and cross excited by opposite end LOs and COs. This is not visible in Figure 9(a) due to the (iii), but in Figure 9(b), one can clearly see the self excitation and the cross excitation by LOs and COs in the red anti-diagonal.
@@
    In-Spread Limit Orders excite other events very similarly between small and large tick stocks.
@@
    Market Orders naturally only occur at the top of the book, however we can see very similar patterns across all types of stocks for Market Orders.
This stylized fact is unique in its ability to provide evidence for the endogeneity of events in the LOB. Furthermore its qualitative uniformity across all stocks tested is especially quite interesting. In the following section we will use this stylized fact as motivation to build a general model for the LOB across all relative tick-sizes.
3 A Wide-Queue Hawkes Model for Limit Order Books
In this section, we utilize the information from the study of stylized facts above to propose a simple extension to the Compound Hawkes Process model proposed in [jain2023hawkes] to account for the nature of both small and large tick stocks.